OpenStudy (anonymous):

Find the equation of the line that is parallel to the line - 2y+7x=3 passes through the point (1,-3). Can you please show details thank you for your help.

OpenStudy (anonymous):

Show the given equation in slope-intercept form y = mx + b. 7/2x -3/2 = y the slope is the coefficient of x or 7/2 A parallel line will have the same slope of 7/2, which is m in the equation. Now you can substitute in to the equation using the point given (1, -3) y = mx + b -3 = 7/2(1) + b solve for b -3 - 7/2 = b -6.5 = b you can write b as a fraction, -13/2 So you have: y = 7/2x - 13/2. To put into standard form, multiply everything by 2 2(y = 7/2x - 13/2) = 2y = 7x-13 13 = 7x - 2y or 7x - 2y = 13

OpenStudy (anonymous):

I am horrible a math, can you tell me how you got 7/2x -3/2 = y the slope is the coefficient of x or 7/2, please. Thank you.

OpenStudy (anonymous):

Oh i got it -2y+7x=3 -2y=3-7x y=(3-7x)/2

OpenStudy (anonymous):

thank you for your help DSS your awesome :)